# Cyber Defense Reinvented: Large Language Models as Threat Intelligence Copilots

### v1

@@ 5.5 Remediation
RF13: CyLens is effective across all remediation tasks. Consistent with RF4 and RF10, we demonstrate CyLens’s effectiveness in handling various remediation objectives as shown in Table 6, including both identification-based (suggesting tools and advisories) and generation-based (producing code patches or mitigation strategies) approaches. In contrast, baseline models, lacking CTI-specific expertise, often exhibit imbalanced performance across remediation tasks. For example, Gemini performs well in generating code patches but struggles to identify the correct advisories.
Figure 7: Evaluation results across multiple CTI tasks based on different levels of organizational interests: (a)-(d) product-level, (e)-(h) vendor-level, and (i)-(l) threat category-level. The evaluation metrics used are accuracy for threat actor attribution, F1 score for attack infrastructure contextualization, accuracy for CVSS base severity classification, and accuracy for third-party advisory suggestion. Model names are SL: SevenLLM, O1: ChatGPT-O1, GM: Gemini-Pro, C1/C8: CyLens-1B/8B.
@@
RF14: Large-scale expertise is crucial. Effective remediation often relies heavily on existing cybersecurity practices [43]. Therefore, embedding large-scale historical remediation knowledge is crucial when developing LLMs for CTI applications. Even for zero-day threats, well-trained LLMs can leverage their inherent knowledge to derive useful insights, improving response effectiveness even in cases where no explicit remediation has been documented.
5.6 Organizational Customization
Previous experiments demonstrate CyLens’s capability in analyzing a wide range of cyber threats, providing a strong foundation for addressing specific organizational needs. Given that threats of particular interest—such as those associated with specific products, vendors, or threat categories—may lack sufficient volume for fine-tuning [58, 35], it is essential to first develop LLMs with comprehensive CTI analysis capabilities. These LLMs can then be incrementally customized to align with specific organizational needs.
We categorize organizational needs into three levels:
@@ 5.6 Organizational Customization
    • Product-Level – Focusing on cyber threats affecting specific products, a common approach among security teams in large enterprises managing product-specific CTI [21].
    • Vendor-Level – Targeting threats that impact an entire vendor (i.e., company), a perspective often adopted by small and medium-sized businesses or industry groups [67].
@@
    • Category-Level –Concentrating on specific threat categories, which is particularly relevant to research institutions conducting specialized cybersecurity studies [49].
Settings. To cover diverse organizational needs, we collect threats associated with (1) specific products (Acrobat, Chrome, or Fedora), (2) specific vendors (openSUSE, GitLab, or Dell), and (3) specific threat categories (Overflow, DoS, or XSS). Thus, we collect a total of nine groups of threats, with each group comprising more than 1,000 threat reports. We allocate 50% of these threat reports—sourced from third-party references in the MITRE-CVE [15], NVD, and Exploit-DB [62] websites [59]—for independently fine-tuning CyLens-1B/8B, while reserving the remaining for evaluation. We compare CyLens to the SOTA model, SevenLLM, and two industry-leading LLMs, GPT-o1 and Gemini. Figure 7 presents the results across multiple CTI tasks.
@@
RF15: CyLens is adaptable to diverse organizational needs. From Figure 7, we observe that CyLens maintains high effectiveness even with independent customization (fine-tuning) across different groups of cyber threats. For instance, CyLens-8B achieves nearly 100% accuracy in providing remediation advisories for various vendors and threat categories. This result further demonstrates its adaptability despite the diverse nature of threats across products, vendors, and categories—factors that significantly impact baseline LLMs, causing their performance to fluctuate. The finding implies CyLens ’s superior capability to generalize while maintaining high effectiveness in tailored needs.
6 Discussion
6.1 Ablation Study
@@
We conduct ablation studies to assess the impact of different components within CyLens. Table 7 presents the results, where we independently exclude pre-training (PT), fine-tuning (FT), and four optional modules: topic modeling (TOM), named entity recognition (NER), relation extraction (REL), and retrieval-augmented generation (RAG). Notably, reasoning (REA) and summarization (SUM) cannot be excluded, as they are essential steps in text generation.
@@
Table 7: Ablation study on CyLens-8B by selecting four representative CTI tasks: (1) attributing threat actors (measured by accuracy), (2) contextualizing attack infrastructure (measured by F1 score), (3) classifying CVSS base severity levels (measured by accuracy), and (4) suggesting advisories for remediation (measured by accuracy).
@@
We observe that pre-training has the most significant impact, as it embeds massive CTI knowledgeinto the model. While fine-tuning is more lightweight, its influence remains non-negligible in refining CyLens ’s CTI capabilities. Additionally, the impact of different modules varies across tasks. For instance, classifying CVSS severity levels does not heavily depend on NER and REL, so removing these modules has minimal effect. However, tasks such as threat actor attribution or advisory suggestion rely on accurately identifying key objects (e.g., malware or third-party vendors), making NER and REL critical for performance. These findings reinforce the necessity of incorporating all proposed techniques to achieve an effective and well-rounded system design for CTI applications.
6.2 Adapting to Computational Constraints
Similar to Meta’s Llama series, which offers models at different scales [56, 57, 92], CyLens is available in three sizes: 1B, 8B, and 70B, allowing it to accommodate varying computational constraints. While CyLens-1B is less effective compared to the 8B and 70B versions, its lightweight design is crucial for resource-limited deployments.
@@ 6.2 Adapting to Computational Constraints
6.3 Privacy-Preserving Deployment
A key advantage of customizing local LLMs is enhanced privacy preservation. Unlike API-based services such as OpenAI’s GPT series, which inherently risk exposing user intentions and confidential data, local models ensure that sensitive information remains under organizational control. API-based solutions may not only transmit proprietary data to external servers but could also use it for model training and further development, raising concerns about data security and regulatory compliance [65, 66, 80, 99, 113].
By developing localized, task-specific LLMs, organizations gain full control over model behavior, customization, and data security. Such privacy-preserving deployment is particularly critical in high-stakes fields like CTI, where threat intelligence, incident reports, and security strategies must remain confidential to prevent adversarial exploitation and ensure compliance with data protection regulations.
6.4 Beyond Specific Needs
Real-world cyber threats are rarely confined to specific products, vendors, or categories [77]. Instead, they manifest within a complex and interconnected ecosystem where vulnerabilities, exploits, and attack methodologies transcend traditional boundaries. This interdependence underscores the necessity of developing CTI solutions that go beyond addressing specific organizational needs (e.g., tailored threat intelligence) and are also capable of capturing the broader, multifaceted threat landscape.
7 Conclusion
This paper introduced CyLens, a LLM-powered copilot for cyber threat intelligence. CyLens streamlines the entire threat management lifecycle, encompassing tasks from threat attribution to remediation. It assimilates knowledge from a vast number of threat reports and utilizes a retrieval module for real-time information, enabling it to adeptly handle both historical and zero-day threats. Rigorous evaluations have demonstrated that CyLens consistently outperforms industry-leading LLMs and SOTA agents across multiple CTI tasks. Its adaptability shines through its customization capabilities, which can be tailored to specific threat categories or organizational requirements. Additionally, the release of its training datasets fosters further research in the field.
@@ Appendix A Pre-Training: Complementary Information
Table 8: The data sources we utilize and the corresponding evidence metadata collected from each database.
@@
Table 9: Base LLM assignment and settings of CyLens’s development, including pre-training and fine-tuning. “Cp” – Corpus; “Tk” – token.
CyLens Base LLM Pre-Training
# Cp # Tk $T$ $T_{1}$ $T_{2}$ $\beta$
CyLens-1B Llama-3.2-1B 28K 0.27B 4 3 3 1.0
CyLens-8B Llama-3.1-8B 115K 1.1B 10 5 5 1.0
CyLens-70B Llama-3.3-70B 271K 2.5B 20 10 10 1.0
A.1 Data Source and Metadata Collection
Complementary to section 3.1, Table 8 provides a detailed overview of the cyber threat databases we utilize, along with the corresponding metadata collected for various CTI tasks (❶-❻) and analytical targets.
MITRE-CVE and NVD. Our collected data is primarily centered around CVEs, leveraging the MITRE-CVE [14] and NVD [59] databases, which serve as well-structured repositories of publicly disclosed vulnerabilities. These databases provide foundational information, including CVE identifiers, descriptions, severity scores, affected products, and references to external sources.
Exploit-DB. To further enrich our dataset and enhance contextual understanding, we incorporate additional cyber threat intelligence (CTI) sources, including Exploit-DB [62]. Exploit-DB provides reports on security advisories and threat intelligence insights, complementing the structured data from MITRE-CVE and NVD. Notably, Exploit-DB is also CVE-centric; however, unlike MITRE-CVE and NVD, it lacks structured connections to additional evidence. Therefore, our data collection primarily focuses on extracting raw threat descriptions and attack vector classifications (e.g., "Local") to supplement our dataset with additional contextual details.
Other MITRE Databases. Beyond these primary databases, we integrate metadata from additional sources to enrich our dataset with diverse perspectives on cyber threats. Specifically, we leverage CWE [16] to systematically categorize software weaknesses, providing a structured foundation for understanding vulnerability root causes. CAPEC [17] offers a comprehensive taxonomy of attack patterns, enabling us to analyze how adversaries exploit known weaknesses in real-world scenarios. MITRE ATT&CK [18] serves as a crucial resource for mapping adversary tactics, techniques, and procedures (TTPs), facilitating in-depth attribution and behavioral analysis.
Third-Party References. Beyond structured frameworks, we incorporate third-party references to capture real-time threat intelligence and contextualized security insights. Among them, GitHub commits to serve as a valuable source for tracking vulnerability disclosures, exploit development, and security patch discussions. Apache security reports [5] provide detailed documentation on vulnerabilities and remediation efforts specific to widely used web and enterprise software. Meanwhile, IBM X-Force threat intelligence [39] contributes curated threat reports, real-world attack observations, and insights into evolving threat landscapes.
@@ A.1 Data Source and Metadata Collection
A.2 Threat Corpus Generation Details
Complementary to section 3.1, we detail the prompts used for generating and refining the threat corpus, specifically in Prompt 1 and 2. During output sampling, we adjust the temperature and top-p parameters within a range of 0.2 to 1.0 to balance creativity and determinism. An example of a generated threat report is provided in Example 4. Additionally, we present the statistical overview of the corpus used for pre-training in Table 9.
Prompt 1
@@ A.2 Threat Corpus Generation Details
Appendix B Fine-Tuning: Complementary Information
Complementary to section 4.1, we present the statistics of the generated fine-tuning dataset in Table 10. Notably, when generating each data instance using a cascading reasoning layout, not all analytical targets are included due to variations in their availability.
Table 10: Statistics of the fine-tuning dataset with a cascading reasoning layout, where each analytical target appears in varying volumes of data. Additionally, different scales of fine-tuning data are used for different CyLens models to align with their respective capacities and optimization needs.
@@ Appendix B Fine-Tuning: Complementary Information
Threat Actor 37.6K 0.8K 11.5K 37.6K
TTPs 35.5K
Exploit Path 21.8K
Affected System 37.6K
Attack Infra 28.3K
Attack Impact 17.6K
CVE Identifier 37.6K
CWE Identifier 37.6K
CVE Identifier 22.9K
CVSS Metrics 37.6K
EPSS Records 33.1K
Tool Use 15.3K
Code Patch 8.2K
Methodology 32.4K
Advisory 27.5K
@@
Furthermore, CyLens-1B, CyLens-8B, and CyLens-70B utilize different volumes of fine-tuning data, reflecting their respective model capacities and optimization requirements. For fine-tuning CyLens-1B, we select 0.8K samples that contain all analytical targets. In contrast, for CyLens-8B and CyLens-70B, we use a progressively larger dataset, with CyLens-8B utilizing a partial subset and CyLens-70B leveraging the full dataset, despite the potential absence of certain analytical targets in some instances.
Appendix C Evaluation: Complementary Information
C.1 Evaluation Data

### v2

@@ 5.5. Remediation
RF12: Broad expertise is crucial. Consistent with RF1 and RF9, we demonstrate CyLens’s effectiveness in addressing a wide range of remediation objectives, as shown in Table 8. These include both identification-based tasks (e.g., suggesting appropriate tools and advisories) and generation-based tasks (e.g., producing code patches or mitigation strategies). In contrast, baseline models often exhibit imbalanced performance across remediation tasks. For instance, while Gemini-Pro performs well in generating code patches, it struggles to correctly identify relevant advisories.
@@
We attribute this disparity to the fact that effective remediation often depends on familiarity with established cybersecurity practices (Jacobs et al., 2020). As such, embedding large-scale historical remediation knowledge is essential when developing LLMs for CTI applications. Notably, even for zero-day threats, a well-trained model like CyLens can leverage its inherent domain knowledge to derive meaningful insights, thereby enhancing response effectiveness in scenarios where no explicit remediation guidance exists.
6. Evaluation Results: Specialized CTI
Despite CyLens’s scalability in capturing a broad threat landscape, real-world CTI often involves targeted interests (Mvula et al., 2024; Optiv, 2025). In this section, we demonstrate CyLens’s adaptability in addressing three specialized groups of threats:
@@ 6. Evaluation Results: Specialized CTI
    • Product – Focusing on cyber threats affecting specific products, a common scenario among security teams in large enterprises managing product-specific CTI (Cybersecurity and Infrastructure Security Agency, 2023).
    • Vendor – Targeting threats that impact an entire vendor (i.e., company), which is a common practice among industry groups, particularly those have focused products or services.(Optiv, 2025).
@@
    • Category – Concentrating on specific threat categories, which is particularly relevant to research institutions conducting specialized cybersecurity studies (Kim and Kwon, 2022).
Specialized Instruction Tuning. Following the same re-training process described in section 3.3, CyLens is independently fine-tuned on threats associated with specific products (MacOS, Chrome, or Linux Kernel), vendors (IBM, Adobe, or Dell), and categories (Overflow, DoS, or XSS). These selections are based on their exploitation popularity as recorded in the MITRE-CVE database (Corporation, 1999). The fine-tuning datasets are derived from a subset of the original collections (Table 11) that include only historical threats, with no overlap with the evaluation dataset (Table 4).
Evaluation Targets. As detailed in section 4, we evaluate CyLens using real-world threat reports collected from multiple platforms (Oracle Corporation, 2025; Red Hat, Inc., 2025; Adobe Systems Incorporated, 2025; Apache Software Foundation, 2025a; Debian Project, 2025; openSUSE Project, 2025; Fedora Project, 2025). From these sources, we extract threats targeting the selected products, vendors, and categories. The corresponding dataset statistics are reported in Table 4.
We compare the specialized CyLens-8B model with two SOTA, SevenLLM and Lily-Cyber, and two industry-leading LLMs, GPT-o1 and Gemini. Additionally, we include a non-specialized CyLens-8B fine-tuned on a broad range of CTI tasks under the same settings as in section 5. The evaluation results across multiple CTI tasks are summarized in Figure 7 and 8. We have the following findings:
@@
RF13: CyLens is adaptable to diverse specialized needs. From Figure 7, we observe that CyLens maintains high effectiveness even with independent instruction-tuning across different groups of cyber threats. For instance, CyLens-8B achieves nearly 100% accuracy in providing remediation advisories for Adobe. This result further demonstrates CyLens’s adaptability despite the diverse nature of threats across products, vendors, and categories—factors that significantly impact baseline LLMs, causing their performance to fluctuate.
RF14: Data scarcity matters. Cyber threats targeting specific products or even entire vendors often suffer from data scarcity, particularly a lack of historical threats that are important for experience-based remediation. This data scarcity can hinder the development of countermeasures, especially for zero-day threats. For example:
Case Study 6.
As of 2024, only 9 cyber threats related to Open Redirection vulnerabilities had been reported for Dell. However, within the first three months of 2025, 2 new threats emerged—targeting Dell Unity (CVE-2025-24381) and Networker (CVE-2025-21104)—highlighting the challenge of generating timely and effective mitigation strategies with limited historical data.
RF15: A broad view of the threat landscape is necessary. When developing CTI solutions for specific products or vendors, relying solely on threat data limited to those entities is often insufficient (as highlighted in RF14). It is essential to consider the broader threat landscape to uncover relevant analytical patterns or remediation strategies. This is demonstrated by the effectiveness of CyLens —both CyLens-1B and CyLens-8B achieve significantly higher accuracy in advisory suggestion for Dell-related threats (¿70%) compared to all baseline methods (¡50%). A representative case study is:
Case Study 7.
In studying the remediation of CVE-2025-21104, CyLens-8B successfully identifies that reconfiguring the web server to enforce strict parsing rules and ignore unexpected Host header values can mitigate the vulnerability—an approach typically incorporated in newer system versions. This insight aligns with Dell’s official advisory DSA-2025-124, which includes a security update for the Dell NetWorker Management Console addressing the issue.
7. Discussion
7.1. Ablation Study
@@
We conduct ablation studies to assess the impact of different components within CyLens. Table 9 presents the results, where we independently exclude re-training (RT), fine-tuning (FT), and four functional modules: TOM, NER, REL, and RAG. Notably, REA and SUM cannot be excluded, as they are essential inference steps.
@@
Table 9. Ablation study on CyLens-8B by selecting four representative CTI tasks: (1) attributing threat actors (measured by accuracy), (2) contextualizing attack infrastructure (measured by F1 score), (3) classifying CVSS base severity levels (measured by accuracy), and (4) suggesting advisories for remediation (measured by accuracy).
@@
We observe that re-training has the most significant impact, as it embeds massive CTI knowledgeinto the model. While fine-tuning is more lightweight, its influence remains non-negligible in refining CyLens ’s CTI capabilities. Additionally, the impact of different modules varies across tasks. For instance, classifying CVSS severity levels does not heavily depend on NER and REL, so removing these modules has minimal effect. However, tasks such as threat actor attribution or advisory suggestion rely on accurately identifying key objects (e.g., malware or third-party vendors), making NER and REL critical for performance. These findings reinforce the necessity of incorporating all proposed techniques to achieve an effective and well-rounded system design for CTI applications.
7.2. Adapting to Computational Constraints
Similar to Meta’s Llama series, which offers models at different scales (Meta-llama, 2024a, b; Touvron et al., 2023), CyLens is available in three sizes: 1B, 8B, and 70B, allowing it to accommodate varying computational constraints. While CyLens-1B is less effective compared to the 8B and 70B, its lightweight design is crucial for resource-limited deployments.
@@ 7.2. Adapting to Computational Constraints
8. Conclusion
We introduce CyLens, a LLM-powered agentic CTI system. CyLens supports the entire threat management lifecycle, encompassing tasks from threat attribution to remediation. It accumulates CTI knowledge from massive threat reports and utilizes a functional module for real-time analysis, enabling it to adeptly handle both historical and zero-day threats. Intensive evaluations have demonstrated the scalability and adaptability of CyLens across massive CTI tasks and tailored needs of threat management. By consistently outperforming industry-leading LLMs and SOTA agents, CyLens offers scalable and adaptive solution and serves as a blueprint for integrating agentic LLMs into CTI practice.
@@ Appendix A Pre-Training: Complementary Information
Table 10. The data sources we utilize and the corresponding evidence metadata collected from each database.
@@
A.1. Data Source and Metadata Collection
Complementary to section 3.2, Table 10 provides a detailed overview of the cyber threat databases we utilize, along with the corresponding metadata collected for various CTI tasks (❶-❺) and analytical targets.
MITRE-CVE and NVD. Our collected data is primarily centered around CVEs, leveraging the MITRE-CVE (Corporation, nd) and NVD (National Institute of Standards and Technology, 2025) databases, which serve as well-structured repositories of publicly disclosed vulnerabilities. These databases provide foundational information, including CVE identifiers, descriptions, severity scores, affected products, and references to external sources.
Exploit-DB. To further enrich our dataset and enhance contextual understanding, we incorporate additional cyber threat intelligence (CTI) sources, including Exploit-DB (Offensive Security, 2025). Exploit-DB provides reports on security advisories and threat intelligence insights, complementing the structured data from MITRE-CVE and NVD. Notably, Exploit-DB is also CVE-centric; however, unlike MITRE-CVE and NVD, it lacks structured connections to additional evidence. Therefore, our data collection primarily focuses on extracting raw threat descriptions and attack vector classifications (e.g., ”Local”) to supplement our dataset with additional contextual details.
Other MITRE Databases. Beyond these primary databases, we integrate metadata from additional sources to enrich our dataset with diverse perspectives on cyber threats. Specifically, we leverage CWE (Corporation, 2006) to systematically categorize software weaknesses, providing a structured foundation for understanding vulnerability root causes. CAPEC (Corporation, 2007) offers a comprehensive taxonomy of attack patterns, enabling us to analyze how adversaries exploit known weaknesses in real-world scenarios. MITRE ATT&CK (Corporation, 2013) serves as a crucial resource for mapping adversary tactics, techniques, and procedures (TTPs), facilitating in-depth attribution and behavioral analysis.
Third-Party References. Beyond structured frameworks, we incorporate third-party references to capture real-time threat intelligence and contextualized security insights. Among them, GitHub commits to serve as a valuable source for tracking vulnerability disclosures, exploit development, and security patch discussions. Apache security reports (Apache Software Foundation, 2025b) provide detailed documentation on vulnerabilities and remediation efforts specific to widely used web and enterprise software. Meanwhile, IBM X-Force threat intelligence (IBM X-Force, 2025) contributes curated threat reports, real-world attack observations, and insights into evolving threat landscapes.
@@ A.1. Data Source and Metadata Collection
A.2. Threat Corpus Generation Details
Complementary to section 3.2, we detail the prompts used for generating and refining the threat corpus, specifically in Prompt 1 and 2. During output sampling, we adjust the temperature and top-p parameters within a range of 0.2 to 1.0 to balance creativity and determinism. An example of a generated threat report is provided in Example A.1. Additionally, we present the statistical overview of the corpus used for pre-training in Table 1.
Prompt 1.
@@ A.2. Threat Corpus Generation Details
Appendix B Instruction Tuning: Complementary Detail
Complementary to section 3.4, we present the statistics of the generated instruction-tuning dataset in Table 11. Notably, when generating each data instance using a cascading reasoning layout, not all analytical targets are included due to variations in their availability.
Table 11. Statistics of the instruction-tuning dataset with a cascading reasoning layout, where each analytical target appears in varying volumes of data. Additionally, different scales of data are used for different CyLens models to align with their respective capacities and optimization needs.
@@ Appendix B Instruction Tuning: Complementary Detail
Threat Actor 37.6K 1.2K 11.5K 37.6K
TTPs 35.5K 1.2K 11.3K 35.5 K
Campaign 21.8K 0.8K 7.9K 21.8K
Affected System 37.6K 1.2K 11.5K 37.6K
Attack Infra 28.3K 0.9K 10.2K 28.3K
Attack Impact 17.6K 0.6K 5.4K 17.6K
CVE Identifier 37.6K 1.2K 11.5K 37.6K
CWE Identifier 37.6K 1.2K 11.5K 37.6K
CVSS Metrics 37.6K 1.2K 11.5K 37.6K
EPSS Records 33.1K 1.3K 12.6K 33.1K
Tool Use 15.3K 0.7K 4.8K 15.3K
Code Patch 8.2K 1.1K 3.1K 8.2K
Methodology 32.4K 1.0K 14.8K 32.4K
Advisory 27.5K 0.8K 9.6K 27.5K
@@
Furthermore, CyLens-1B, CyLens-8B, and CyLens-70B utilize different volumes of fine-tuning data, reflecting their respective model capacities and optimization requirements. For fine-tuning CyLens-1B, we select a smaller volume of samples that contain all analytical targets. In contrast, for CyLens-8B and CyLens-70B, we use a progressively larger dataset, with CyLens-8B utilizing a partial subset and CyLens-70B leveraging the full dataset, despite the potential absence of certain analytical targets in some instances.
Appendix C Inference Module: Complementary Detail
This section detail the implementation of six NLP modules as described in section 3.5.
C.1. Topic Modeling
To perform topic modeling, we leverage BERTopic, a SOTA topic modeling framework that integrates transformer-based embeddings with clustering and class-based TF-IDF. In our implementation, we design a two-stage prompt-based workflow to integrate BERTopic with CyLens for dynamic topic modeling. In the first stage, CyLens is prompted to assess whether topic modeling is needed based on the input text’s characteristics. CyLens receives instructions:
Prompt 3.
Necessity of Topic Modeling.
Given the input query, determine whether topic modeling is appropriate. Hint: topic modeling is suitable when the texts cover multiple themes or the user seeks to explore underlying topics.
CyLens evaluates the semantic content and returns a binary decision along with reasoning. Once CyLens determines that topic modeling is appropriate, it proceeds to construct a structured call to BERTopic via a second-stage prompt. This prompt guides CyLens to assemble the necessary inputs for BERTopic and invoke it accordingly. The prompt is:
Prompt 4.
Calling Topic Modeling.
You have identified that topic modeling is required. Use BERTopic to extract topics from the following documents. For each topic, return a descriptive label, a list of representative keywords, and top exemplar documents.
CyLens structures this as a function call or prompt completion depending on the system design. BERTopic is then executed with the provided documents, and the results are passed back to CyLens, which can further head to the next NLP module.
C.2. Named Entity Recognition (NER)
To perform NER, we adopt a prompt-based inference approach in CyLens in place of traditional sequence labeling methods. Rather than fine-tuning, CyLens is queried at runtime with carefully crafted prompts to identify and categorize named entities directly from the input text. The prompt is
Prompt 5.
NER Prompt (An Example for Attribution Task).
System Prompt: You are a cybersecurity threat intelligence assistant specialized in named entity recognition (NER). Your task is to extract and categorize entities relevant to threat attribution from unstructured text. Focus on identifying entities that help answer the question ”Who is responsible for the attack?” and ”How was the attack carried out?” Instructions: Given a cybersecurity-related document or report excerpt, extract all relevant named entities and classify them into the following categories: Threat Actor: Names or aliases of individuals or groups suspected or known to be behind the activity. $\cdots$ Return the results in the following structured JSON format.
CyLens then returns a list of entities, each annotated with its type and surface form.
C.3. Relation Extraction
To extract semantic relationships between entities in cybersecurity texts, we implement relation extraction through prompt-based inference in CyLens. This method allows us to bypass traditional supervised training by leveraging the CyLens ’s few-shot and zero-shot reasoning capabilities. At inference time, we provide the CyLens with structured prompts that instruct it to identify and describe relationships between extracted entities, such as connections between threat actors and malware, or between attack campaigns and their target sectors. A typical prompt format is
Prompt 6.
Relation Extraction Prompt
System Prompt: You are a cybersecurity threat intelligence reasoning assistant. Your task is to perform relation extraction to identify and structure meaningful semantic relationships between entities mentioned in cybersecurity texts, particularly in the context of threat attribution. The goal is to infer who did what to whom, how, when, and using what tools or infrastructure. Instructions: Given a cybersecurity-related passage (such as a threat report or incident summary) and its extracted named entities, identify relevant relationships between entities and represent them as a list of (subject, relation, object) triples. Focus on attribution-relevant relations. Use only clearly supported relationships based on the input. If no relationships are found, return an empty list. Return the results in this structured JSON format.
C.4. Retrieval-Augmented Generation (RAG)
To enhance the quality of CyLens inference, we implement a RAG framework that integrates external web-based APIs during inference. This bridges generative reasoning with real-time access to structured and unstructured data sources, such as open-source intelligence feeds, public CTI repositories, and search APIs.
The RAG workflow begins with the CyLens analyzing the user’s input to determine whether external retrieval is necessary. For example, if the input question involves current threat actor activity, recent campaign data, or infrastructure details (e.g., domains, malware hashes), the system triggers a retrieval phase. In this phase, CyLens is prompted to generate a structured search query that captures the essential information needs of the input. The prompt format follows a declarative schema, such as: “Formulate a concise search query to retrieve current information about [topic], including actors, tools, targets, and known affiliations.” The generated query is sent to a relevant external data source via API—for instance, querying public CTI portals (e.g., MITRE ATT&CK, VirusTotal, Shodan, or threat feed aggregators) or web search APIs (e.g., Bing Search or Google Custom Search).
Upon receiving the retrieved content (typically one or more passages or structured documents), the system constructs a new input prompt to CyLens, embedding the original question and the retrieved evidence into the context window. This second-stage prompt is crafted using templates such as: “Given the following query and retrieved documents, generate a comprehensive, evidence-based answer. Use only the provided content.” This constraint guides CyLens to ground its reasoning in the externally retrieved information, thereby improving factual accuracy and traceability.
C.5. Reasoning
To support advanced analytical capabilities in CyLens, we implement modular reasoning and summarization components using prompt engineering. Rather than relying solely on fine-tuned models for specific downstream tasks, our architecture employs LLMs (i.e., CyLens itself) as general-purpose reasoning engines, guided by carefully constructed prompts. This allows CyLens to remain flexible, easily extensible, and capable of performing domain-specific tasks such as cyber threat correlation, threat actor profiling, and intelligence synthesis without retraining.
The reasoning module is designed to interpret and connect extracted entities, attributes, and contextual cues across threat intelligence documents. We use prompt engineering to simulate multi-step analytical thinking, instructing the LLM to infer relationships, fill in missing links, and assess causality. For instance, prompts are constructed as: “Given the following entities and attack context, identify the most plausible attribution and justify your reasoning step-by-step.” This encourages the model to break down the logic behind associations, such as linking a threat actor to a campaign based on observed TTPs, tool reuse, or geopolitical motivation. We also use chain-of-thought prompting and structured templates (e.g., tabular reasoning) to improve traceability and consistency of the model’s analytical conclusions.
C.6. Summarization
The summarization module in CyLens is responsible for generating high-level, human-readable briefings from dense technical threat reports or multi-source intelligence inputs. To achieve this, we implement domain-adapted summarization prompts that guide CyLens to preserve critical cybersecurity-relevant information such as named entities, TTPs, indicators of compromise (IOCs), and timeline markers. The prompt format typically follows: “Summarize the following document into a concise threat intelligence report, highlighting threat actors, attack vectors, affected sectors, and key dates.” Depending on the use case, we support multiple summarization styles, including executive summaries, technical briefs, and IOC-centric extracts. We also integrate few-shot examples when needed to adapt the tone or structure to a specific user role (e.g., SOC analyst vs. CISO).
Appendix D Evaluation: Complementary Information
Figure 8. Additional evaluation results across on attack infrastructure contextualization (F1 score) and CVSS base severity classification (accuracy), complemented to Figure 7.
D.1. Evaluation Data